Suppose the supply curve for wool is given by Q s = P, where Q s is the quantity offered for sale when the price is P. Also suppose the demand curve for wool is given by Q d = 10 − P + I , where Q d is the quantity of wool demanded when the price is P and the level of income is I. Assume I is an exogenous variable.

Question:Find the equilibrium price and quantity if I is unknown (I is an exogenous variable & we are looking for a parametric equilibrium). For what values of I will the equilibrium exist? If I increases by a little bit, how will equilibrium price and quantity behave i.e. find the derivative of equilibrium price and quantity with respect to I ?
Answer #1

| Qs = P Qd = 10 - Pt I At equilibrium, asa ad or p = 10-p+I a, 2P = 10-1 cop 5 - 7 - So, g- 5- I quantity whone I is- CPR Q